# MERA: A Comprehensive LLM Evaluation in Russian

### v2

@@ Appendix B Motivation for the Selection of the Number of Few-shot Examples
Appendix C Baseline Details
C.1 Random Baseline Details
@@
  • For each task from the ruEthics dataset, we use random.randint from Python to provide each of the five outputs as an answer.
  • For each task from the ruModAr dataset and from the ruMultiAr dataset we use random.randint from Python to provide integer in range [$-10^{6}$; $10^{6}$] as an answer.
@@
C.2 Model Baseline Details
For all models we set up dtype=auto to ensure correct precision used and use batch size of one for better reproducibility323232lm-evaluation-harness issue 704: “For some models and prompts, the loglikelihood changes with the batch size”.
For decoder models, we use hf-causal-experimental and for encoder-decoder models, we use hf-seq2seq internal model class type of customized lm-harness code.
@@ C.2 Model Baseline Details
C.3 Task-specific Details
@@
  • PARus, ruOpenBookQA, MultiQ, CheGeKa were taken from RSG and TAPE with no changes and, therefore, we report the baselines of the originals research Shavrina et al. (2020b); Taktasheva et al. (2022).
  • USE human baseline is based on the official examination statistics. ruHumanEval includes specific tasks that a regular annotator cannot solve due to a lack of programming skills. These tasks have straightforward algorithmic solutions, so we assign each pass@k metric the value of 1 (the metrics value in Table 5 is multiplied by 100).
C.4 Annotation Procedure Details
@@
Annotation procedure on Toloka:
@@
The annotation procedure via the ABC platform slightly differs. The quality monitoring there is performed by moderators, while the other annotation steps remain the same as for the Toloka annotation procedure.
Tab. 7 summarizes all general details concerning the human evaluation for each project. The dataset size for ruModAr, ruMultiAr, BPS, LCS and SimpleAr datasets differs as the samples for annotation have been randomly chosen from test sets following the uniform distribution. The tasks from these datasets are guaranteed to have a single correct answer that can found using a strict algorithm, so there is no need in larger amount of samples to estimate human performance on such tasks.
Task name Total Item Pay rate Dataset size
Toloka
MathLogicQA
\$233.9
\$0.041
\$1.03/hr
1143
RCB
\$73.46
\$0.034
\$2.61/hr
438
ruModAr
\$190.08
\$0.021
\$1.23/hr
1800
ruMultiAr
\$75.94
\$0.025
\$1.01/hr
600
LCS
\$14.5
\$0.029
\$1.73/hr
100
BPS
\$10.17
\$0.02
\$3.2/hr
100
ruWorldTree
\$81.31
\$0.031
\$2.36/hr
525
RWSD
\$27.05
\$0.021
\$1.48/hr
260
ruMMLU
\$192.38
\$0.04
\$1.58/hr
961
SimpleAr
\$28.98
\$0.029
\$3.33/hr
200
ruHateSpeech
\$40.42
\$0.031
\$3.22/hr
265
ruHHH
\$170.08
\$0.026
\$1.92/hr
178
ruDetox
\$364.11
\$0.03
\$3.83/hr
800
ABC
ruTiE
\$27.18
\$0.063
\$0.707/hr
430
ruEthics
\$175.22
\$0.091
\$1.77/hr
1935
Table 7: The details of human baseline evaluation. Total is the budget spent to annotate the tasks employed for metric evaluation. Item is the weighted average reward of the annotator for one item. Pay rate is the hourly rate computed as a simple average of pay rates based on time spent annotating one row and the reward for this row. Dataset size refers to the total number of documents in the test part of the dataset.

### v3

@@ Appendix D Motivation for the Selection of the Number of Few-shot Examples
Appendix E Baseline Details
E.1 Random Baseline Details
@@
  • For each task from the ruEthics dataset, we sample a random integer from a range of [0, 1] for each of the five labels using random.randint.
  • For each task from the ruModAr dataset and from the ruMultiAr dataset we sample a random integer from a range of [$-10^{6}$; $10^{6}$] as an answer using random.randint.
@@
E.2 Model Baseline Details
We run all models on NVIDIA A100 GPUs [^34] https://www.nvidia.com/en-us/data-center/a100 with torch 2.0.0 Paszke et al. (2019) and transformers 4.36.2 Wolf et al. (2020).
For all models we set up dtype=auto to ensure correct precision used and use batch size of one for better reproducibility353535lm-evaluation-harness issue 704: “For some models and prompts, the log-likelihood changes with the batch size”.
For decoder models, we use hf-causal-experimental and for encoder-decoder models, we use hf-seq2seq internal model class type of customized lm-eval code.
@@ E.2 Model Baseline Details
For the davinci-002 model, we used openai==1.10.0 version. The scoring took place on 25 Jan 2024, which may be necessary for the reproducibility of the results.
E.3 Human Baseline Details
@@
  • PARus, ruOpenBookQA, MultiQ, CheGeKa were taken from RSG Shavrina et al. (2020b) and TAPE Taktasheva et al. (2022) with no changes and, therefore, we report the baselines of the original research.
  • USE human baseline is based on the official examination statistics [^36] https://doc.fipi.ru/ege/analiticheskie-i-metodicheskie-materialy/2019/russkiy_yazyk_2019.pdf.
  • ruHumanEval includes specific tasks that a regular annotator cannot solve due to a lack of programming skills. These tasks have straightforward algorithmic solutions, so we assign each pass@k metric the value of 1 (the value of the metric in Tab. 5 is multiplied by 100).
Appendix F Annotation Procedure Details
@@
The new datasets were created from scratch, but their design process differed. Some were generated through the proposed methodology based on English counterparts (e.g., ruModAr, SimpleAr, ruMultiAr). Several datasets were created manually by various experts without the crowdsource platform usage (e.g., ruHumanEval, ruHHH, ruTiE, ruMMLU test parts). The remaining datasets were created using crowdsourced platforms ABC or Toloka (e.g., MathLogicQA, ruHateSpeech, ruEthics, ruMMLU public test split). Details for the latter can be found in Tab. 7.
The human baseline was also obtained using Toloka and ABC platforms. We use the following annotation procedure on Toloka for a human baseline:
@@
The annotation procedure via the ABC platform slightly differs. The quality monitoring on the platform is performed by moderators, while the other annotation steps remain the same as for the Toloka annotation procedure.
Tab. 8 summarizes all general details concerning the human evaluation for each project.
It should be noted that the example number for ruModAr, ruMultiAr, BPS, LCS, and SimpleAr datasets differs from the size of the original test as the samples for annotation have been randomly chosen from test sets following the uniform distribution. The tasks from these datasets are guaranteed to have a single correct answer that can be found using a strict algorithm, so there is no need for a larger amount of samples to estimate human performance on such tasks.
Task name Total Item Pay rate Example number Overlap IAA
MathLogicQA \$586.28 \$0.046 \$1.24/hr 2570 3 89%
ruHateSpeech \$4082.57 \$0.037 \$2.32/hr 20479 3 87%
ruEthics \$45.59 \$0.3 \$6.84/hr 152 3 N/A*
ruMMLUpublic test \$7770 \$0.098 \$1.97/hr 15858 5 81%
Table 7: The details of datasets collection and verification. Total is the budget spent to annotate the tasks employed for metric evaluation. Item is the weighted average reward of the annotator for one item. Pay rate is the hourly rate computed as a simple average of pay rates based on time spent annotating one row and the reward for this row. Example number refers to the total number of samples processed while collecting or verifying the dataset. Overlap is the median number of votes per dataset sample averaged across all annotation tasks for the same dataset (if more than 1 task provided). IAA stands for inter-annotator agreement, which is the share of the answer voted for by the most annotators among all answers averaged across all dataset samples and all annotation tasks for the same dataset (if more than 1 task provided). *Not available for ruEthics as the annotators’ answers are barely comparable since each actant may be described by different word combinations from the texts.
Task name Total Item Pay rate Example number Overlap IAA
Toloka MathLogicQA \$233.9 \$0.041 \$1.03/hr 1143 5 93%
RCB \$73.46 \$0.034 \$2.61/hr 438 4 57%
ruModAr \$190.08 \$0.021 \$1.23/hr 1800 5 95%
ruMultiAr \$75.94 \$0.025 \$1.01/hr 600 5 95%
LCS \$14.5 \$0.029 \$1.73/hr 100 5 46%
BPS \$10.17 \$0.02 \$3.2/hr 100 5 95%
ruWorldTree \$81.31 \$0.031 \$2.36/hr 525 5 88%
RWSD \$27.05 \$0.021 \$1.48/hr 260 5 80%
ruMMLUtest \$192.38 \$0.04 \$1.58/hr 961 5 76%
SimpleAr \$28.98 \$0.029 \$3.33/hr 200 5 98%
ruHateSpeech \$40.42 \$0.031 \$3.22/hr 265 5 94%
ruHHH \$70.55 \$0.019 \$3.28/hr 178 5 77%
ruDetox \$364.11 \$0.03 \$3.83/hr 800 4 N/A*
ABC ruTiE \$27.4 \$0.064 \$0.713/hr 430 5 90%
ruEthics \$175.22 \$0.091 \$1.77/hr 1935 5 N/A*
Table 8: The details of human baseline evaluation. Total is the budget spent to annotate the tasks employed for metric evaluation. Item is the weighted average reward of the annotator for one item. Pay rate is the hourly rate computed as a simple average of pay rates based on time spent annotating one row and the reward for this row. Example number refers to the total number of samples used for human baseline evaluation. Overlap is the median number of votes per dataset sample. IAA stands for inter-annotator agreement, which is the share of correct answers among all answers averaged across all dataset samples. *Not available for ruEthics as there are no target variables, for ruDetox due to annotating the already existing detoxified texts.